A 48.0 kg in-line skater begins with a speed of 2.2 m/s. Friction also does (-150) J of work on her. Assume that se did not push on the ground any more. If her final speed is 5.9 m/s,


a) determine the change (final-initial) in her gravitational potential energy.


b) By how much, and in which direction (up or down), has her weight changed?


Please answer using grade 11 Conservation of Energy and Momentum .



1
Expert's answer
2020-05-08T16:25:52-0400

As per the given question,

mass of the line skater "(m)=48 kg"

speed of the line skater "(u)=2.2 m\/sec"

work done by friction force "(w)=-150J"

her final speed "(v)=5.0m\/sec"

a) Applying the conservation of energy,

Initial potential energy "(U_1)=\dfrac{mu^2}{2}"

Final energy "(U_2)=\dfrac{mv^2}{2}-150"

so, change in PE"=
\dfrac{48\times 5.9^2}{2}-\dfrac{48\times 2.2^2}{2}-150=(835.44-232.32-150 )J=453.12J"


b)Weight of the skater is working downwards, because as she is moving downwards, her speed is getting increase, it means it is getting accelerate towards the ground.